# X-Boundary: Establishing Exact Safety Boundary to Shield LLMs from Multi-Turn Jailbreaks without Compromising Usability

### v1

@@ Abstract
Despite the rapid development of safety alignment techniques for LLMs, defending against multi-turn jailbreaks is still a challenging task. In this paper, we conduct a comprehensive comparison, revealing that some existing defense methods can improve the robustness of LLMs against multi-turn jailbreaks but compromise usability, i.e., reducing general capabilities or causing the over-refusal problem. From the perspective of mechanism interpretability of LLMs, we discover that these methods fail to establish a boundary that exactly distinguishes safe and harmful feature representations. Therefore, boundary-safe representations close to harmful representations are inevitably disrupted, leading to a decline in usability. To address this issue, we propose X-Boundary to push harmful representations away from boundary-safe representations and obtain an exact distinction boundary. In this way, harmful representations can be precisely erased without disrupting safe ones. Experimental results show that X-Boundary achieves state-of-the-art defense performance against multi-turn jailbreaks, while reducing the over-refusal rate by about 20% and maintaining nearly complete general capability. Furthermore, we theoretically prove and empirically verify that X-Boundary can accelerate the convergence process during training.
Warning: this paper includes examples that may be offensive or harmful.
large language model, safety defense
@@ 1 Introduction
In this paper, we pioneeringly adapt and comprehensively compare four single-turn defense methods against multi-turn jailbreaks, including supervised fine-tuning (SFT) (Ren et al., 2024b), Direct Preference Optimization (DPO) (Jiang et al., 2024), Gradient Ascent (GA) (Zhang et al., 2024b), and circuit breaking (CB) (Zou et al., 2024). Although these methods can reduce the attack success rate (ASR) of multi-turn jailbreaks, they compromise the usability of LLMs, i.e., reducing general capability or leading to severe over-refusal issues. For example, the coding capability relatively reduces about 20% after SFT and the over-refusal rate exceeds 50% after GA [^1] Tested on on Qwen2.5-7B-Chat and Mistral-7B-Instruct-v0.2, respectively. Please see Table 1 for detailed results.. None of them successfully achieve a balance between defense robustness and usability. Over-refusal means that LLMs may unreasonably reject safe prompts (Röttger et al., 2023; Shi et al., 2024; Cui et al., 2024). For instance, Fig. 1 shows that LLMs refuse to answer a safe prompt “What are the harmful effects of meth on the human body?”.
@@ 2.1 Adaption and Evaluation of Single-Turn Defense Methods Against Multi-Turn Jailbreaks
Evaluation of defense robustness. To evaluate the robustness of these defense methods, we test them against three SOTA multi-turn jailbreak attacks: ActorAttack (Ren et al., 2024b), RedQueen (Jiang et al., 2024) and Crescendo (Russinovich et al., 2024). The evaluation metric used is the Attack Success Rate (ASR), defined as the proportion of attack attempts that successfully elicit harmful content from the LLMs. A lower ASR indicates greater robustness against multi-turn jailbreak attacks.
@@ Appendix A Additional Results
A.1 The Trade-Off between Robustness and General Capability
@@
A.2 Ablation Study on Llama-3-8B-Instruct and Mistral-7B-Instruct-v0.2
@@
A.3 Effects of the Size of Boundary-Safe Data
@@
A.4 Effects of Separate Loss and Boundary-Safe Data on the Representation Distribution
@@
A.5 Details about Representation Visualization
@@ Appendix D Case Study
Cases of Over-Safety In Fig. 13 and Fig. 14, we present two examples from OKTest and OR-Bench, comparing the responses of Llama-3-8B-Instruct with different safety strategies. X-Boundary provides accurate and useful responses to boundary-safe queries, e.g., “Is there an effective program in place to educate children about the dangers of molestation? ”, unlike previous methods which incorrectly treat them as harmful queries and refuse to respond. This lies in X-Boundary effectively increases the distance between harmful representations and boundary-safe representations, allowing for a clear distinction between harmful queries and boundary-safe queries.

### v2

@@ Abstract
Despite the rapid development of safety alignment techniques for LLMs, defending against multi-turn jailbreaks is still a challenging task. In this paper, we conduct a comprehensive comparison, revealing that some existing defense methods can improve the robustness of LLMs against multi-turn jailbreaks but compromise usability, i.e., reducing general capabilities or causing the over-refusal problem. From the perspective of mechanism interpretability of LLMs, we discover that these methods fail to establish a boundary that exactly distinguishes safe and harmful feature representations. Therefore, boundary-safe representations close to harmful representations are inevitably disrupted, leading to a decline in usability. To address this issue, we propose X-Boundary to push harmful representations away from boundary-safe representations and obtain an exact distinction boundary. In this way, harmful representations can be precisely erased without disrupting safe ones. Experimental results show that X-Boundary achieves state-of-the-art defense performance against multi-turn jailbreaks, while reducing the over-refusal rate by about 20% and maintaining nearly complete general capability. Furthermore, we theoretically prove and empirically verify that X-Boundary can accelerate the convergence process during training. Please see our code at: https://github.com/AI45Lab/X-Boundary. Warning: this paper includes examples that may be offensive or harmful.
Xiaoya Lu1,2,*Dongrui Liu1,*Yi Yu2Luxin Xu3Jing Shao2,†
1 School of Electronic Information and Electric Engineering, Shanghai Jiao Tong University
2 Shanghai Artificial Intelligence Laboratory
3 University of Electronic Science and Technology of China
footnotetext: Correspondence to: Jing Shao <shaojing@pjlab.org.cn>
@@ 1 Introduction
In this paper, we pioneeringly adapt and comprehensively compare four single-turn defense methods against multi-turn jailbreaks, including supervised fine-tuning (SFT) (Ren et al., 2024b), Direct Preference Optimization (DPO) (Jiang et al., 2024), Gradient Ascent (GA) (Zhang et al., 2024b), and circuit breaking (CB) (Zou et al., 2024). Although these methods can reduce the attack success rate (ASR) of multi-turn jailbreaks, they compromise the usability of LLMs, i.e., reducing general capability or leading to severe over-refusal issues. For example, the coding capability relatively reduces about 20% after SFT and the over-refusal rate exceeds 50% after GA. None of them successfully achieve a balance between defense robustness and usability. Over-refusal means that LLMs may unreasonably reject safe prompts (Röttger et al., 2023; Shi et al., 2024; Cui et al., 2024). For instance, Fig. 1 shows that LLMs refuse to answer a safe prompt “What are the harmful effects of meth on the human body?”.
@@ 2.1 Adaption and Evaluation of Single-Turn Defense Methods Against Multi-Turn Jailbreaks
Evaluation of defense robustness. To evaluate the robustness of these defense methods, we test them against three SOTA multi-turn jailbreak attacks: ActorAttack (Ren et al., 2024b), RedQueen (Jiang et al., 2024) and Crescendo (Russinovich et al., 2024). The evaluation metric used is the Attack Success Rate (ASR), defined as the proportion of attack attempts that successfully elicit harmful content from the LLMs. A lower ASR indicates greater robustness against multi-turn jailbreak attacks. We also evaluate the compliance rates of LLMs with harmful requests in HarmBench and the results are listed in the Appendix A.1.
@@ 4.2 Main Results
Table 3: Comparison of existing defense methods and X-Boundary on reasoning models.
Models Methods Single & Multi-Turn ASR (%) $\downarrow$ Over-Refusal Rate (%) $\downarrow$ General Capability (%) $\uparrow$
HarmBench ActorAttack RedQueen XSTest OKTest OR-Bench PHTest AIME GPQA LiveCode
DeepSeek- R1-Distill- Llama-8B Vanilla 55.42 83.50 58.50 0.40 1.67 1.00 4.67 50.00 50.00 40.00
SFT 16.25 44.50 28.50 58.40 34.67 32.67 65.00 40.00 44.95 35.10
DPO 17.50 67.00 36.50 0.40 1.67 1.33 7.33 50.00 46.97 42.40
CB 52.80 77.50 55.00 1.60 3.33 3.67 8.00 46.67 46.97 40.65
X-Boundary 13.33 34.00 27.00 2.40 3.00 7.33 12.00 50.00 50.00 39.86
DeepSeek- R1-Distill- Qwen-7B Vanilla 60.42 79.00 60.50 1.60 2.00 1.33 2.67 53.30 48.99 39.76
SFT 48.33 59.50 42.50 7.60 9.00 11.00 12.67 46.67 48.99 36.44
DPO 52.92 74.00 48.50 1.60 3.67 3.00 8.67 53.33 50.00 40.08
CB 57.92 54.00 55.50 5.33 4.00 5.67 9.00 46.67 46.97 39.33
X-Boundary 20.80 43.50 36.00 2.40 3.00 7.33 12.00 53.33 50.51 40.02
@@
4.3 Defense Performance on Reasoning Models
Recently, Jiang et al. (2025); Zhou et al. (2025) have highlighted significant safety risks in large reasoning models, such as DeepSeek-R1 (Guo et al., 2025). Notably, the reasoning outputs of R1 models pose even greater safety concerns than their final answers. In this subsection, we explore the implementation of X-Boundary to enhance the safety defenses of reasoning models. As shown in Table 3, X-Boundary effectively reduces the ASR of both single-turn and multi-turn attacks, while not causing severe over-refusal issues or a decline in reasoning ability in distilled reasoning models. Surprisingly, the CB appears to be less effective for reasoning models, possibly because of differences in their representation space compared to standard models. According to the theoretical insights in Section 3.3, X-Boundary reduces the difficulty of training and accelerates the training process. Therefore, X-Boundary has an advantage over CB. The effect of defense methods on the reasoning ability of R1 models is discussed in Appendix A.2.
@@ Appendix A Additional Results
A.1 The compliance rates in HarmBench
We evaluate the basic safety level of the LLMs using HarmBench (Mazeika et al., 2024). Table 5 shows both existing defense methods and X-Boundary can effectively reduce the compliance rates to harmful requests in HarmBench without using attack methods.
Table 5: The compliance rates to harmful requests in HarmBench after using existing defense methods and X-Boundary.
Methods Llama-3- 8B-Instruct Qwen2.5- 7B-Chat Mistral-7B- Instruct-v0.2 Qwen2.5- 14B-Chat
Vanilla 11.67 26.25 56.67 15.83
SFT 1.25 5.42 7.08 1.25
DPO 0.83 2.92 13.33 5.83
GA 5.00 3.75 9.17 /
CB 1.67 4.58 13.33 4.58
X-Boundary 1.25 4.58 13.33 2.92
A.2 The effect of defense methods on the LLMs’ reasoning ability
Large reasoning models often rely on generating lengthy reasoning paths for inference. Therefore, we conducted a statistical analysis of the output length of large reasoning models employing various defense mechanisms. As shown in Table 6, while X-Boundary does not lead to a degradation in general capability, it results in shorter output lengths, which may indirectly impact reasoning performance. Exploring strategies to prevent the reduction in output length represents a promising direction for future research.
Table 6: Comparison of pass@1 accuracy and average output token length across different defense methods on reasoning model
Models Methods AIME2024 GPQA LiveCode
pass@1 Length (Avg.) pass@1 Length (Avg.) pass@1 Length (Avg.)
DeepSeek- R1-Distill- Llama-8B Vanilla 50.00 15672.07 50.00 8910.93 40.00 6457.43
SFT 44.95 13678.53 40.00 8699.93 35.10 6804.28
DPO 46.97 15716.27 50.00 8489.33 42.40 6301.96
CB 46.97 15488.23 46.97 9088.78 40.65 6479.9
X-Boundary 50.00 13310.90 50.00 8233.20 39.86 6498.04
DeepSeek- R1-Distill- Qwen-7B Vanilla 53.33 11046.63 48.99 8592.54 39.76 6683.22
SFT 46.67 13844.87 48.99 8176.29 36.44 6825.17
DPO 53.33 12063.57 50.00 8344.05 40.08 6694.74
CB 46.97 12609.93 46.97 8356.40 39.33 6536.76
X-Boundary 53.33 12959.73 50.51 8237.67 40.02 6583.29
A.3 The Trade-Off between Robustness and General Capability
@@
A.4 Ablation Study on Llama-3-8B-Instruct and Mistral-7B-Instruct-v0.2
@@
A.5 Effects of the Size of Boundary-Safe Data
@@
A.6 Effects of Separate Loss and Boundary-Safe Data on the Representation Distribution
@@
A.7 Details about Representation Visualization
@@ Appendix D Case Study
Cases of Over-Refusal In Fig. 13 and Fig. 14, we present two examples from OKTest and OR-Bench, comparing the responses of Llama-3-8B-Instruct with different safety strategies. X-Boundary provides accurate and useful responses to boundary-safe queries, e.g., “Is there an effective program in place to educate children about the dangers of molestation? ”, unlike previous methods which incorrectly treat them as harmful queries and refuse to respond. This lies in X-Boundary effectively increases the distance between harmful representations and boundary-safe representations, allowing for a clear distinction between harmful queries and boundary-safe queries.
@@
Cases of Safety Defense on Reasoning Model In Fig. 18, with X-Boundary, dangerous content generated in the reasoning process is automatically converted into gibberish. Additionally, we use carefully designed rules to detect gibberish, allowing us to terminate the thought process early and replace the response with a refusal answer.
@@
Figure 18: Safety defense example on reasoning models.